# Attractor Stability of Boolean networks under noise

### v1

@@ Abstract
We study the impact of noise on attractor dynamics in Boolean networks, focusing on their stability and transition behaviors. By constructing attractor matrices based on single-node perturbations, we propose a framework to quantify attractor stability and identify dominant attractors. We find that attractors are more stable than predicted by basin sizes, showing the importance of dynamical structure in noisy environments. In addition, under global perturbations, basin sizes dictate long-term behavior; under local noise, however, attractor dominance is determined by noise-induced transition patterns rather than basin sizes. Our results show that transition dynamics induced by stochastic perturbations provide an efficient and quantitative description for the attractor stability and dynamics in Boolean networks under noise.
@@ I Introduction
Boolean networks have been widely used as mathematical models for various biological systems, in particular the study of gene regulatory networks [1, 2, 3], metabolic networks [4], signal transduction networks [5], and neural networks [6]. These models provide a simple yet powerful framework to capture the underlying interactions and dependencies between components [7]. In the classical Boolean networks, each node represents a binary state (00 or $1$), and the state of each node is updated based on a Boolean function that depends on the states of its neighboring nodes [1, 8, 9]. Many variants of Boolean network models have since been proposed to better reflect biological complexity, and a wide range of dynamical behaviors have been extensively studied [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. One of the key features of Boolean networks is that their state trajectories eventually converge to stable configurations (fixed points) or recurring patterns (limit cycles), so called “attractors” [1, 20, 21]. Attractors play a crucial role in understanding the long-term behavior of the system, as they correspond to different functional states in biological systems [1, 11].
@@
Noise in biological systems is ubiquitous, arising from intrinsic fluctuations in biochemical reactions and external environmental variations [22, 23, 24]. Noise can randomly alter the states of biological systems, independent to the system’s underlying rules and thus plays a significant role in shaping dynamics [22, 24, 25, 26]. In Boolean networks, noise can be implemented as random state flips of individual nodes, that are independent of the Boolean update rules [25, 27, 26, 28]. From the perspective of attractor dynamics, such noise can induce significant changes: rather than remaining permanently trapped in a single attractor, the system may transition between different attractors over time [25, 27, 29]. As a result, understanding how attractors behave under noise is essential for capturing the full dynamical properties of biological systems [29]. It also allows for a more complete view of biological stability and variability under noisy environments. A related modeling framework is the probabilistic Boolean networks [3, 30], where state transitions are governed by probabilities rather than deterministic rules, sharing similar notions of stochastic effects in Boolean networks. In contrast to the probabilistic framework, where randomness is incorporated into the update rules, the noise considered here operates independently of the system’s deterministic rules.
@@
Given the importance of understanding how noise influences attractor dynamics, it is crucial to analyze the stability of attractors under noise and the long-term dynamics of Boolean networks. Among existing approaches, Derrida analysis has been widely used to capture the spread of perturbations by assessing the average sensitivity in Boolean networks [20, 31, 32, 33]. However, this method is limited in its ability to address noise-driven transitions between attractors, as it does not explicitly include stochastic effects and, moreover, operates at the node level rather than the attractor level [28]. To this end, we propose a framework for computing transition probabilities between attractors under noise. This framework enables us to systematically assess the stability of attractors in Boolean networks and understand how noise impacts their long-term dynamics.
@@
The remainder of this paper is organized as follows. In Sec. II, we begin by introducing Boolean networks under noise. Next, in Sec. III, we describe the mathematical framework for quantifying transition probabilities between attractors and present our approach for constructing transition matrices to analyze the transition patterns of attractors. In Sec. IV, we analyze how noise alters the long-term behavior of Boolean networks by quantifying attractor stability, identifying dominant attractors, and characterizing the structure of noise-induced transitions. We also apply our analysis to real-world Boolean networks to examine the relevance of our findings in biological systems. Finally, in Sec. V, we summarize our findings and discuss the implications of our results and potential directions for future research.
@@ II Boolean networks under noise
where $F$ is a Boolean function that governs the dynamics of nodes based on the states of its input nodes at time $t$. The Boolean function $F$ is predefined and remains fixed throughout the evolution of the system. The system updates all nodes in parallel at each time step. Since the dynamics are deterministic, the system eventually converges to a single attractor in any finite system. These attractors can be either fixed points where the state remains a single state or limit cycles where the system oscillates among a finite set of states [8].
@@
An example of a Boolean network with $N=3$ is depicted in Fig. 1(a). The wiring structure shows regulatory influences between nodes through directed edges, with Boolean functions specifying how each node updates its state based on its inputs. The corresponding state space of the network is shown in Fig. 1(b), where each node represents a possible configuration of the system and directed edges indicate deterministic evolution between states according to the Boolean update rules. As shown in Fig. 1(b), all trajectories eventually lead to one of the attractors, which can be either a fixed point or a limit cycle.
@@
We introduce stochastic effects by adding noise to the system in the form of random flipping of the state of a node. Under noisy conditions, the system evolves through a combination of deterministic Boolean rules and stochastic perturbations. When noise occurs, the state of an affected node is randomly altered, leading to deviations from the purely deterministic trajectory [25, 34, 35, 26]. Such fluctuations can drive the system to transition between attractors, preventing it from remaining in a single attractor indefinitely. In this study, we assume that noise occurs relatively infrequently. If noise were too frequent, the system would no longer follow the logic of a Boolean network but instead behave like a randomly fluctuating system, which would lose biological relevance.
Our main interest in this model lies in how noise affects attractor dynamics in Boolean networks. To be specific, we examine the stability of individual attractors by analyzing how likely the system is to remain within the same attractor after random perturbations. We also explore which attractors tend to dominate the steady-state behavior of the system under noise, by characterizing the transition patterns between attractors and quantifying the probabilities of noise-induced transitions.
III Transition matrices of Boolean networks under noise
@@
In this section, we propose a framework for constructing a transition matrix that encodes the probabilities of transitions between attractors due to noise. Using this matrix, we analyze the stability of attractors and the consequences of transitions between them in noisy environments.
@@ III.1 Transition probabilities between attractors
The process for computing the transition matrix under noise is as follows. We first simulate the time evolution of states in a given Boolean network and identify all possible attractors. Each state in the network converges to one of these attractors after passing through transient states. We introduce noise by flipping the value of a single randomly chosen node in a given attractor state, and simulate the dynamics to determine the resulting attractor. This process mimics external perturbations that disrupt the deterministic evolution of the system. Since noise acts intermittently, the Boolean network remains in attractors for most of the time rather than in transient states. Therefore, we restrict our analysis to attractor states, as the system stays in them with high probability. By repeating this process for every state in each attractor, we estimate the transition probabilities between attractors.
@@
III.2 Stability and dominance of attractors from transition matrices
From the computed transition probabilities, we construct a transition matrix
@@ III.2 Stability and dominance of attractors from transition matrices
where each element $m_{\alpha\beta}$ represents the probability of transitioning from attractor $\alpha$ to attractor $\beta$ under noise. The matrix $M$ defines a Markov chain over the set of attractors, where transitions are governed by noise-induced perturbations. Each element $m_{\alpha\beta}$ represents the probability of transitioning from attractor $\alpha$ to attractor $\beta$, satisfying $0\leq m_{\alpha\beta}\leq 1$ and $\sum_{\beta}m_{\alpha\beta}=1$ for all $\alpha$.
The diagonal elements $m_{\alpha\alpha}$ represent the probability that the system remains in the same attractor $\alpha$ after flipping a node. We interpret this quantity as a measure of attractor stability; that is, higher values indicate more robust attractors, whereas lower values imply a higher tendency to transition to other attractors under noise. The off-diagonal elements $m_{\alpha\beta}$ with $\alpha\neq\beta$ quantify the probabilities of noise-induced transitions between different attractors. Thus, the transition matrix $M$ captures the overall stability landscape of the attractor structure in the presence of noise.
In addition to stability, $M$ also provides insight into the dominance of attractors through its long-term stationary distribution. The transition matrix has a principal eigenvalue $\lambda_{1}=1$ with a corresponding eigenvector $\vec{v}$ satisfying $M\vec{v}=\vec{v}$. The components of $\vec{v}$ form the stationary distribution over attractors, where $v_{\alpha}$ represents the fraction of time the system spends in attractor $\alpha$. In this sense, attractors with larger $v_{\alpha}$ are considered dominant, as the system tends to stay in them more frequently under noise.
@@ III.3 Global randomization and attractor basins
To set a baseline for comparison, we consider a global perturbation scenario in which noise simultaneously and randomly alters the states of all nodes. In this case, the entire state rather than a single node undergoes a random change. Specifically, each node in the Boolean network is independently assigned 00 or $1$ with equal probability when noise is applied. We refer to this scenario as “global randomization”, as flipping is applied to the entire network rather than to individual nodes. It is important to note that this type of perturbation is fundamentally different from the local noise model considered in our study, where only a single node’s state is flipped. The global randomization erases all memory of the previous state since the system is entirely reset to a random point in the state space.
In this setting, the transition matrix is determined solely by the basin sizes of the attractors. Here, the “basin” of an attractor refers to the set of all initial states that converge to that attractor, and its relative size, $b$, reflects how likely a randomly selected initial state converges to that attractor. Since global randomization selects states uniformly at random from the full state space, the probability $m_{\alpha\beta}$ that global noise drives the system from attractor $\alpha$ to attractor $\beta$ depends solely on the basin size $b_{\beta}$ of $\beta$. To be specific, $m_{\alpha\beta}=b_{\beta}/(N\ell_{\alpha})$, where $b_{\beta}$ denotes the normalized basin size of attractor $\beta$. As a result, all columns of the matrix $M$ are identical, and the eigenvector associated with the principal eigenvalue is given by the basin size distribution, $b_{\alpha}$.
Therefore, under global noise, attractors with larger basins exhibit higher probabilities of being reached, and basin size becomes the primary indicator of long-term behavior in this scenario. This observation provides a mathematical rationale for the use of basin size as an indicator of attractor importance or stability, as larger basins are generally associated with greater robustness to perturbations [36, 21]. However, this correspondence does not hold under the local noise model considered in our study, where transitions depend on how individual attractor states respond to single-node perturbations. Nevertheless, basin size serves as a baseline for evaluating how local noise modifies attractor dynamics.
@@ IV Results
In this section, we analyze the stability of attractors and the probabilities of transitions between them. We then study how noise influences attractor dynamics in random Boolean networks to understand fundamental properties. To this end, we generate random regular networks with degree $k$ as the underlying structures, and assign Boolean functions randomly according to a given bias parameter $p$, which represents the probability that the output of a Boolean function is 00. For comparison across different values of $k$, we use the average sensitivity $s=2p(1-p)k$ [20, 31] as a unified control parameter. The sensitivity $s$ quantifies the expected number of node state changes caused by flipping a single input and serves as a standard measure of the system’s dynamical sensitivity. We also apply our analysis to empirical networks to explore the real-world implications.
IV.1 Stability of attractors under noise
In this section, we study the stability of attractors in Boolean networks by analyzing the system’s ability to return to the same attractor after a perturbation caused by local noise. In this sense, the diagonal elements $m_{\alpha\alpha}$ of the transition matrix represent the stability of attractor $\alpha$. To assess attractor stability at the network level, we compute the trace of the transition matrix, $\mathrm{Tr}(M)$. A larger trace indicates that attractors in Boolean networks are more stable to noise. As a baseline for comparison, we consider global randomization in which the system’s state is reset entirely at random. In this null model, the trace of the transition matrix equals one by definition, as it results from the normalization of basin sizes. We therefore use $\mathrm{Tr}(M)=1$ as a reference point, and evaluate the stability of Boolean networks by measuring their trace values.
Our results in Fig. 2(a) show that the trace $\mathrm{Tr}(M)$ is consistently greater than unity across various values of the connectivity $k$ and sensitivity $s$, on random regular networks with $N=20$. These observations suggest that Boolean networks have an intrinsic resilience to local perturbations that is higher than expected from basin size. While basin size captures the probability of reaching an attractor under global randomization, it is insufficient to account for how the system tends to remain within the same attractor under local noise. Our findings suggest the importance of assessing stability through dynamical measures, $\mathrm{Tr}(M)$, which more directly reflect the system’s resilience to noise.
Figure 2: (a) The trace $\mathrm{Tr}(M)$ of transition matrix for random Boolean networks on random regular networks with size $N=20$ and various degrees $k$ with respect to the sensitivity $s$ is shown. (b) The trace normalized by the number of attractors, $\mathrm{Tr}(M)/n_{\mathrm{att}}$ is shown. Inset shows the average number $n_{\mathrm{att}}$ of attractors for given $k$ and $s$.
In Fig. 2(b), we show the average attractor stability, defined as $\mathrm{Tr}(M)/n_{\mathrm{att}}$, where $n_{\mathrm{att}}$ is the number of attractors in a given network. This normalization is introduced because the number of attractors varies across different network instances and parameter settings. The results show that average attractor stability tends to decrease as the average sensitivity $s$ increases. Networks with different values of connectivity $k$ but the same value of sensitivity $s$ exhibit similar levels of attractor stability. This indicates that the sensitivity parameter $s$ plays a role as a more reliable predictor of network-level stability than $k$ or $p$ alone.
Figure 3: Similarities $\sigma$ of states within attractors are shown as a function of the sensitivity $s$. Results are obtained from random regular networks with $N=15$, and various values of $k$, averaged over $10^{6}$ independent runs.
To understand the source of high attractor stability, we analyze the internal similarity of states within attractors. Consider attractor $\alpha$, which consists of $\ell_{\alpha}$ binary states $\{\mathbf{x}^{(\alpha)}_{1},\mathbf{x}^{(\alpha)}_{2},\dots,\mathbf{x}^{(\alpha)}_{\ell_{\alpha}}\}$, where $\ell_{\alpha}$ is the length of the attractor $\alpha$, i.e., the number of distinct states in the attractor. Each state $\mathbf{x}^{(\alpha)}_{i}$ is a binary vector of length $N$. The similarity $\sigma^{\alpha}$ for attractor $\alpha$ is defined as the average probability that two randomly selected states in attractor $\alpha$ have the same value at a given position. Let $n_{1}^{(\alpha,i)}$ and $n_{0}^{(\alpha,i)}$ respectively be the number of ones and zeros in attractor $\alpha$ in a position $i$. Then the similarity $\sigma$ can be computed as
$\sigma_{\alpha}=\frac{1}{N}\sum_{i=1}^{N}\frac{\binom{n_{0}^{(\alpha,i)}}{2}+\binom{n_{1}^{(\alpha,i)}}{2}}{\binom{\ell_{\alpha}}{2}}.$ (4)
For fixed-point attractors with length $\ell_{\alpha}=1$, we define the similarity $\sigma_{\alpha}=1$, since all positions are trivially identical. This allows a unified similarity measure for all attractors regardless of their lengths.
To compute the average similarity $\sigma$ across all attractors in the network, we introduce the following approximation, instead of explicitly computing all binomial coefficients over attractor state. Let $\mu_{i,\alpha}=n_{1}^{(\alpha,i)}/\ell_{\alpha}$ be the fraction of states with value $1$ at position $i$. When the attractor length $\ell_{\alpha}$ is large, we approximate the overall similarity of attractors as
$\sigma\approx\frac{1}{n_{\mathrm{att}}}\sum_{\alpha=1}^{n_{\mathrm{att}}}\left[\frac{1}{N}\sum_{i=1}^{N}\left(\mu_{i,\alpha}^{2}+\left(1-\mu_{i,\alpha}\right)^{2}\right)\right].$ (5)
In this approximation, the similarity $\sigma$ represents the probability that two randomly selected elements are equal, either both 00 with probability $(1-\mu)^{2}$ or both $1$ with probability $\mu^{2}$.
Figure 3 shows the similarity $\sigma$ of states in an attractor as a function of average sensitivity $s$, for various values of connectivity $k$. If attractors were composed of randomly distributed states, $\sigma$ becomes $1/2$. However, the observed values are consistently higher than $1/2$, indicating that the states within each attractor tend to be more similar to one another than randomly chosen states. Therefore, local perturbations, i.e., flipping a single node are less likely to push the system out of its attractor and it partially explains why Boolean networks exhibit higher attractor stability under noise. In addition, the pattern of $\sigma$ in Fig.3 shows a decreasing trend with increasing $s$. This pattern shares a similarity with that of the average attractor stability $\mathrm{Tr}(M)/n_{\mathrm{att}}$ shown in Fig.2(b). It suggests that $\sigma$ provides a useful qualitative indicator of attractor stability in Boolean networks.
IV.2 Dominance and stationary distribution of attractors

### v2

# Dynamics of attractor transitions in Boolean networks under noise

@@ Abstract
Biological systems operate under persistent noise, which can alter system states and induce transitions between attractors. Here, we study the attractor dynamics of Boolean networks focusing on the transitions between attractors induced by noise. By computing transition probabilities between attractors, we present methods at the attractor level to determine dominance, stability, and diversity of attractors, and systematically compare local and global noise. Whereas global noise leads to attractor behavior dictated primarily by basin sizes, local noise produces structured transition patterns characterized by enhanced stability, non-trivial dominance patterns, and broader exploration of the attractor space. Our work offers insight into the dynamics of attractors, showing the importance of transition patterns under noise.
Biological systems are inherently noisy, with stochastic fluctuations often driving transitions between functional states. Despite this, much of the analysis of Boolean networks, one of the common theoretical frameworks for gene regulation and other biological processes, has overlooked the dynamic impact of noise. Motivated by this gap, this study provides a framework to address the attractor dynamics in Boolean networks under noise. By constructing and analyzing transition probabilities among attractors, we quantify dominance and stability of attractors, and the diversity of attractor distributions. We further compare the effects of local state-flip noise and global randomization, showing the distinct patterns of attractor dynamics.
@@ I Introduction
Boolean networks have been widely used as mathematical models for various biological systems, in particular the study of gene regulatory networks [1, 2, 3], metabolic networks [4], signal transduction networks [5], and neural networks [6]. These models provide a simple yet powerful framework to capture the underlying interactions and dependencies between components [7]. In the classical Boolean networks, each node represents a binary state (0 or $1$), and the state of each node is updated based on a Boolean function that depends on the states of its neighboring nodes [1, 8, 9]. Many variants of Boolean network models have since been proposed to better reflect biological complexity, and a wide range of dynamical behaviors have been extensively studied [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. One of the key features of Boolean networks is that their state trajectories eventually converge to stable configurations (fixed points) or recurring patterns (limit cycles), so called “attractors” [1, 21, 22]. Attractors play a crucial role in understanding the long-term behavior of the system, as they correspond to different functional states in biological systems [1, 11].
@@
Noise in biological systems is ubiquitous, arising from intrinsic fluctuations in biochemical reactions and external environmental variations [23, 24, 25]. Noise can randomly alter the states of biological systems, independent to the system’s underlying rules and thus plays a significant role in shaping dynamics [23, 25, 26, 27]. In Boolean networks, noise can be implemented as random state flips of nodes, that are independent of the Boolean update rules [26, 28, 27, 29]. From the perspective of attractor dynamics, such noise can induce significant changes: rather than remaining permanently trapped in a single attractor, the system may transition between different attractors over time [26, 28, 30]. As a result, understanding how attractors behave under noise is essential for capturing the full dynamical properties of biological systems [30]. It also allows for a more complete view of biological stability and variability under noisy environments. A related modeling framework is the probabilistic Boolean networks [3, 31], where state transitions are governed by probabilities rather than deterministic rules, sharing similar notions of stochastic effects in Boolean networks. In contrast to the probabilistic framework, where randomness is incorporated into the update rules, the noise considered here operates independently of the system’s deterministic rules.
@@
Given the importance of understanding how noise influences attractor dynamics, it is crucial to analyze the transition patterns between attractors under noise and their long-term dynamics in Boolean networks. Among existing approaches, Derrida analysis has been widely used to capture the spread of perturbations by assessing the average sensitivity in Boolean networks [21, 32, 33, 34]. However, this method is limited in its ability to address the the transition patterns of attractors induced by noise, as it does not explicitly include stochastic effects and, moreover, operates at the node level rather than the attractor level [29]. To this end, we propose a framework for analyzing attractor dynamics based on transition probabilities between attractors under noise. This framework enables us to systematically assess dynamical structure of attractors in Boolean networks and understand how noise shapes their long-term behaviors.
@@
The remainder of this paper is organized as follows. In Sec. II, we begin by introducing Boolean networks under noise. Next, in Sec. III, we describe the mathematical framework for quantifying transition probabilities between attractors and for describing the dynamics of Boolean networks on the attractor level. In Sec. IV, we show how our framework can be used to explore attractor-level dynamics in Boolean networks. Our analysis allows us to extract several attractor-level quantities, such as the frequencies with which attractors are visited, attractor stability, and the diversity of attractor distributions. We also apply our analysis to real-world Boolean networks to examine the applicability our methods to empirical data. Finally, in Sec. V, we summarize our results and discuss the implications of our study and potential directions for future research.
@@ II Boolean networks under noise
where $F$ is a Boolean function that governs the dynamics of nodes based on the states of its input nodes. The Boolean function $F$ is predefined and remains fixed throughout the evolution of the system. The system updates all nodes in parallel at each time step. Since the dynamics are deterministic, the system eventually converges to a single attractor in any finite system. These attractors can be either fixed points where the state remains a single state or limit cycles where the system oscillates among a finite set of states [8].
@@
An example of a Boolean network with $N=3$ is depicted in Fig. 1(a). The wiring structure shows regulatory influences between nodes through directed edges, with Boolean functions specifying how each node updates its state based on its inputs. The corresponding state space of the network is shown in Fig. 1(b), where each node represents a possible configuration of the system and directed edges indicate deterministic evolution between states according to the Boolean update rules. As shown in Fig. 1(b), all trajectories eventually lead to one of the attractors, which can be either a fixed point or a limit cycle. As an example, consider the trajectory starting from state $(1,1,0)$ in Fig. 1(b). The state first moves to $(1,0,0)$ and then settles into the fixed-point attractor $(0,0,0)$. The non-repeating sequence $(1,1,0)\rightarrow(1,0,0)$ constitutes the transient period before the system reaches its attractor.
@@
We introduce stochastic effects by adding noise to the system in the form of random flipping of the state of nodes. Under noisy conditions, the system evolves through a combination of deterministic Boolean rules and stochastic perturbations. When noise occurs, the state of affected nodes is randomly altered, leading to deviations from the purely deterministic trajectory [26, 35, 36, 27]. Such fluctuations can drive the system to transition between attractors, preventing it from remaining in a single attractor indefinitely. In this study, we assume that noise occurs relatively infrequently compared to the typical transient time required for the system to relax back to an attractor after a perturbation. If noise were too frequent, the system would no longer follow the logic of a Boolean network but instead behave like a randomly fluctuating system, which would lose biological relevance.
Our main objective is to propose a framework for attractor-level analysis in Boolean networks under noise. Using this framework, we study various aspects of attractor dynamics, such as how stable each attractor is to perturbations, which attractors are more frequently visited in the presence of noise, and how noise shapes the patterns of transitions between attractors.
III Attractor dynamics via transition probabilities
@@
In this section, we propose a framework for analyzing attractor dynamics under noise by constructing a transition matrix that encodes the probabilities of transitions between attractors. Using this matrix, we examine the transition patterns of attractors in noisy environments, including both local and global noise, and show how it can be used to predict key quantities of attractor dynamics.
@@ III.1 Transition probabilities between attractors
We first simulate the time evolution of states in a given Boolean network and identify all possible attractors. Each state in the network converges to one of these attractors after passing through transient states. We introduce lcoal noise by flipping the value of a randomly chosen node in a given attractor state, and simulate the dynamics to determine the resulting attractor. This process mimics external perturbations that disrupt the deterministic evolution of the system. Since noise acts intermittently, the Boolean network remains in attractors for most of the time rather than in transient states. Therefore, we restrict our analysis to attractor states, as the system stays in them with high probability. By repeating this process for every state in each attractor, we estimate the transition probabilities between attractors.
@@
III.2 Attractor dynamics from transition matrices
From the computed transition probabilities, we can construct a transition matrix
@@ III.2 Attractor dynamics from transition matrices
where each element $m_{\alpha\beta}$ represents the probability of transitioning from attractor $\alpha$ to attractor $\beta$ under noise. The matrix $M$ defines a Markov chain over the set of attractors, where transitions are governed by perturbations. Each element $m_{\alpha\beta}$ represents the probability of transitioning from attractor $\alpha$ to attractor $\beta$, satisfying $0\leq m_{\alpha\beta}\leq 1$ and $\sum_{\beta}m_{\alpha\beta}=1$ for all $\alpha$.
The diagonal elements $m_{\alpha\alpha}$ represent the probability that the system remains in the same attractor $\alpha$ after flipping a node. We interpret this quantity as a measure of attractor stability; that is, higher values indicate more stable attractors, whereas lower values imply a higher tendency to transition to other attractors under noise. The off-diagonal elements $m_{\alpha\beta}$ with $\alpha\neq\beta$ quantify the probabilities of the attractor transitions from $\alpha$ to $\beta$.
The transition matrix $M$ captures the attractor-level dynamics in the Boolean networks in the presence of noise. To be specific, let $\vec{u}(t)=(u_{1}(t),u_{2}(t),\dots,u_{n}(t))$ denote the vector whose component $u_{\alpha}(t)$ is the probability that the system occupies the attractor $\alpha$ at time $t$. The evolution of these probabilities follows
$\vec{u}(t+1)=M^{\mathsf{T}}\vec{u}(t),$ (4)
where $M^{\mathsf{T}}$ denotes the transpose of the matrix $M$. Thus, the transition matrix $M$ encodes the coarse-grained dynamics at the attractor level.
We can also obtain the dominance of attractors through its long-term stationary distribution. The transition matrix has a principal eigenvalue $\lambda_{1}=1$ with a corresponding eigenvector $\vec{v}$ satisfying $M^{\mathsf{T}}\vec{v}=\vec{v}$. The components of $\vec{v}$ form the stationary distribution over attractors, where $v_{\alpha}$ represents the fraction of time the system spends in attractor $\alpha$. In this sense, attractors with larger $v_{\alpha}$ are considered dominant, as the system tends to stay in them more frequently under noise. In our analysis, we focus only on non-degenerate cases where the principal eigenvalue has a unique eigenvector. In degenerate cases, the stationary distribution is not uniquely defined. Such degeneracy typically arises when the attractor transition structure decomposes into multiple disconnected components, preventing a unique value of dominance. For this reason, degenerate cases are excluded from our analysis.
@@ III.3 Global randomization and attractor basins
For comparison, we consider a global noise scenario in which noise simultaneously and randomly alters the states of all nodes. In this case, the entire state rather than a single node undergoes a random change. Specifically, each node in the Boolean network is independently assigned 0 or $1$ with equal probability when noise is applied. We refer to this scenario as global randomization, as flipping is applied to the entire network rather than to individual nodes. It is important to note that this type of perturbation is fundamentally different from the local noise model considered in our study, where only a single node’s state is flipped. The global randomization erases all memory of the previous state since the system is entirely reset to a random point in the state space.
In this setting, the transition matrix is determined solely by the basin sizes of the attractors. Here, the “basin” of an attractor refers to the set of all initial states that converge to that attractor, and its relative size, $b$, reflects how likely a randomly selected initial state converges to that attractor. Since global randomization selects states uniformly at random from the full state space, the probability $g_{\alpha\beta}$ that global noise drives the system from attractor $\alpha$ to attractor $\beta$ depends solely on the basin size $b_{\beta}$ of $\beta$. To be specific, $g_{\alpha\beta}=b_{\beta}/N$, where $b_{\beta}$ denotes the normalized basin size of attractor $\beta$. As a result, all columns are identical, and the eigenvector associated with the principal eigenvalue is given by the basin size distribution, $b_{\alpha}$.
Therefore, under global noise, attractors with larger basins exhibit higher probabilities of being reached, and basin size becomes the primary indicator of long-term behavior in this scenario. This observation provides a mathematical rationale for the use of basin size as larger basins are associated with greater importance of attractors [37, 22, 38, 39, 40]. However, this correspondence does not hold under the local noise model, where transitions depend on how individual attractors respond to single-node perturbations.
@@ IV Results
In this section, we analyze the dynamics of Boolean networks under noise by using the attractor-level description. By representing hopping between attractors by a transition matrix, we can study attractor dynamics in a simple and compact manner. To this end, we generate random regular networks with degree $k$ as the underlying structures, and assign Boolean functions randomly according to a given bias parameter $p$, which represents the probability that the output of a Boolean function is 0. In our analysis, we allow self-regulation. For comparison across different values of $k$, we use the average sensitivity s=2p(1p)ks=2p(1-p)k [21, 32] as a unified control parameter. The sensitivity $s$ quantifies the expected number of node state changes caused by flipping a single input and serves as a standard measure of the system’s dynamical sensitivity. We also apply our analysis to empirical Boolean networks.
IV.1 Dominance and stationary distribution of attractors